OpenStudy (anonymous):

Please help 3 questions! 1. Which is an equation of the line that contains the points (0,3) and (-2,4) A. 2x + y = 3 B. x + 2y = 6 C. 2x + y = 0 D. x - 2y = 6 2. A line contains the point (-3,-1) and has a slope of 1/3. which equation represents this line? A. x + 3y = 0 B. x - 3y = 0 C. 3x + y = 0 D. 3x - y = 0 3. Which is an equation of the line that contains the points (0,2) and (4,0)? A. x + 2y = 4 B. x - 2y = 4 C. 2x + y = 4 D. 2x - y = 4

OpenStudy (shinalcantara):

1.) get the slope of the line.. \[m = \frac{ 4-3 }{ -2-0 }\] you're given with the y-intercept as (0,3) so plug it all in to y = mx+b where m is the slope and b is the y-intercept

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

so it would be -1/2?

OpenStudy (shinalcantara):

yep now you have m=-1/2 and b=3 so substitute

OpenStudy (anonymous):

y=-1/2x+3?

OpenStudy (anonymous):

How do i substitute it?

OpenStudy (anonymous):

i put?

OpenStudy (shinalcantara):

\[y = mx + b\] \[y = -\frac{ 1 }{ 2 }x + 3\] find the LCD for the terms at the right \[y = \frac{ (-1)x+2(3) }{ 2 }\] cross multiplt 2 to the left \[2y = -x + 6\] now rearrange by transposing -x to the left

OpenStudy (anonymous):

do you subtract 6 from both sides?

OpenStudy (anonymous):

oh wait

OpenStudy (shinalcantara):

nope just add x to both sides so that the right side will be left with the constant

OpenStudy (anonymous):

x+2y=6?

OpenStudy (shinalcantara):

yep

OpenStudy (anonymous):

oh

OpenStudy (shinalcantara):

2. since you have a point(-3,-1) and a slope=1/3 just use the slope equation \[m = \frac{ y-y _{1} }{ x-x _{1} }\] where m=1/3 and (x1,y1) is (-3,-1) just substitute

OpenStudy (anonymous):

so it would be m = 1/3 ?/-3,-1

OpenStudy (shinalcantara):

yep \[m = \frac{ y-y _{1} }{ x-x _{1} }\] \[\frac{ 1 }{ 3 } = \frac{ y-(-1) }{ x-(-3) }\]

OpenStudy (anonymous):

Oh

OpenStudy (shinalcantara):

simplify that

OpenStudy (shinalcantara):

remember that when dealing with these problems, it will just revolve around 2 formulas/equations: Slope-intercept form: \[y=mx+b\] and Slope formula \[m = \frac{ y-y _{1} }{ x-x _{1} }\]

OpenStudy (shinalcantara):

what did you get?

OpenStudy (anonymous):

I don't know how to do O_O

OpenStudy (anonymous):

1/3 = -1y/-3x?

OpenStudy (shinalcantara):

\[\frac{ 1 }{ 3 } = \frac{ y-(-1) }{ x-(-3) }\] \[\frac{ 1 }{ 3 } = \frac{ y+1 }{ x+3 }\] cross multiply \[1(x+3) = 3(y+1)\] distribute 1 and 3 to respective quantities \[1(x)+1(3) = 3(y) +3(1)\] \[x+3 = 3y + 3\] subtract both side with (x+3) \[x+3-x-3 = 3y + 3 - x -3\] \[0 = 3y -x\] \[3y-x=0\]

OpenStudy (anonymous):

So B?

OpenStudy (shinalcantara):

yep

OpenStudy (shinalcantara):

3. you're given with y-intercept as (0,2) so m=2 just find the slope using the slope formula

OpenStudy (anonymous):

so use m = y1-y2/x1-x2?

OpenStudy (anonymous):

or y=mx+b?

OpenStudy (shinalcantara):

the first one m=...

OpenStudy (anonymous):

ok

OpenStudy (shinalcantara):

use the given points (0,2) and (4,0)

OpenStudy (anonymous):

so it would be m = 2-0/0-4?

OpenStudy (shinalcantara):

yep

OpenStudy (anonymous):

which would be m = 2/-4?

OpenStudy (shinalcantara):

yep or reducing it to lowest form m = -1/2

OpenStudy (shinalcantara):

now you are ready to substitute to the equation y=mx+b

OpenStudy (shinalcantara):

m = -1/2, b = 2

OpenStudy (anonymous):

ok so y = 1/2x + 2

OpenStudy (shinalcantara):

\[y = -\frac{ 1 }{ 2 }x + 2\] getting the LCD \[y = \frac{ -1(x) +2(2) }{ 2 }\] cross multiply \[2y = -x + 4\] add x to both right and left

OpenStudy (anonymous):

Uh so x - 2y = 4?

OpenStudy (anonymous):

it would cancel it out and it would be 2x = 4?

OpenStudy (shinalcantara):

2y +x = -x + 4 +x

OpenStudy (shinalcantara):

2y + x = 4

OpenStudy (anonymous):

oh so A?

OpenStudy (anonymous):

Thank you for helping me

OpenStudy (shinalcantara):

yep

OpenStudy (anonymous):

I give you medal and fan you

OpenStudy (shinalcantara):

yw :) just keep on practicing for it to retain on your memory

OpenStudy (anonymous):

OK ty